Circular No. 8651

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005mh–2005mr CBET 339 contains the details for ten supernovae (designated 2005mh–

2005mq) fainter than mag g = 20 reported by J. Frieman on behalf of the Sloan Digital Sky Survey II collaboration; six of the new SDSS supernovae are type-Ia events, one is a probable type-Ia event (2005mh), one a type-II event (2005mk), one to be a probable type-IIn event (2005mj), and one is a type-Ib supernova (2005mn). CBET 340 contains details of a probable type-Ia supernova (mag 24–25), designated 2005mr and apparently in a galaxy with redshift z = 0.68, found by G. R. Meurer and colleagues (the “Probing Evolution And Reionization Spectroscopically” Science Team) in spectroscopic data obtained with the Hubble Space Telescope’s Advanced Camera for Surveys Wide Field Camera.

SUPERNOVA 2005ms IN UGC 4614 T. Puckett and A. Kroes report the discovery of an apparent supernova

(mag 17.9) on an unfiltered CCD frame taken with the Puckett Observatory 0.50-m automated supernova patrol telescope in Ellijay, GA, on Dec. 27.30 UT (and confirmed at mag 17.1 by J. Newton, Portal, AZ, on frames taken on Dec. 29.40 with a 0.35-m reflector). The new object is located at α = 8^{h}49^{m}14^{s}.34, δ = +36^{o}07′47′′.9 (equinox 2000.0), which is 25′′.0 west and 35′′.9 north of the center of UGC 4614. SN 2005ms is not present on frames taken by Puckett on Dec. 11 (limiting mag 19.2)

SUPERNOVAE 2005ev, 2005ew, 2005mt–2005mx Below are details for seven supernovae discovered by the “Nearby Su-

pernova Factory” collaboration on NEAT images taken at Palomar. All are type-Ia supernovae except for 2005mt (type IIn), 2005mw (type II), and 2005mx (type II). (Additional information given on CBETs 244, 344.)

SN  2005 UT  α_{2000}  δ_{2000}  Mag.
2005ev Oct. 6.4 2^{h}31^{m}04.^{s}93 +27^{o}42′06′′.5 16.0
2005ew Oct. 4.5 3 39 23.74 +35 02 49.0 15.1
2005mt Feb. 3.45 10 24 12.9 − 3 44 50.6 18.0
2005mu May 19.5 21 35 52.5 −26 27 03.5 17.4
2005mv June 21.5 20 45 16.8 − 3 49 35.7 19.6
2005mw June 25.4 17 37 44.2 +11 09 03.7 19.2
2005mx July 1.5 21 40 08.2 +10 39 14.0 19.4
2005 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green